OpenStudy (anonymous):

y=-3/7x-12 i need to solve for the x and y axis.

OpenStudy (anonymous):

plug in x=0 to find y int plug in y=0 to find x int

OpenStudy (anonymous):

i dont understand , how would you do it as a fraction though?

OpenStudy (anonymous):

y=-3/7(0)-12 =-12 is the y intercept 0=-3/7x-12 12=-3/7x 12 divided by -3/7=-28 is the x intercept

OpenStudy (anonymous):

i cant figure out how you got -28 for x?

OpenStudy (anonymous):

12 divided by -3/7 would be like 12/1 x -7/3 = -84/3 =-28

OpenStudy (anonymous):

so you flipped -7/3 ?

OpenStudy (anonymous):

yep 8)

OpenStudy (anonymous):

ok , so the answer is -28?

OpenStudy (anonymous):

i believe soo